### v1

@@ 1 Introduction
Autonomous underwater vehicles (AUVs) and remotely operated vehicles (ROVs) have been proposed as scalable alternatives or supplementary systems for diver-based surveys and trawling methods to estimate species abundance, biodiversity, and animal behavior in a variety of environments ranging from coral reefs to deeper benthic communities [clarke_using_2009, powell_first_2022, obura_coral_2019]. Many of these approaches rely on cameras to record videos or imagery during vehicle surveys, which can then be used to visually count animals or study behavior directly.
Recent studies of marine animal interactions with underwater vehicles [kruusmaa_salmon_2020, maxeiner_social_2023] suggest that the presence of vehicles themselves can lead to a bias in the estimation of these quantities, impacting scientists’ ability to trust those measurements. For instance, if animals are repelled or attracted to a vehicle, this will cause under- or over-estimation of their populations, respectively. However, many of these studies were conducted on animals in captivity, which often do not reflect the behaviors of animals in the wild [turko_physiological_2023].
Campbell et al. [campbell_assessment_2021] provides one of the first studies of marine animal-robot interaction in the wild. They used a camera trap setup and measured fish responses to AUVs, ROVs, and Towed Vehicles (TVs). Their initial findings suggest that different vehicles do produce noticeable shifts in fish counts. However, due to the logistical challenges of their setup, it is difficult to reproduce and validate a variety of other hypotheses in other types of environments. The goals of this work is to:
@@
    Provide a theoretical framework to identify and systematically minimize measurement bias of AUVs for collecting data about marine animal behavior.
@@
    Provide and validate simple protocol for collecting data in-situ to inform the model.
@@ 2.1 Model of animal response
As in [campbell_assessment_2021], we hypothesize that certain species of fish show behavioral responses to vehicles, but that it is highly correlated with their sensing modalities. However, many of these quantities are difficult to measure and so, regardless of modality, there is some threshold distance at which their sensing or reaction is limited. It is known that many prey fish will shelter in the coral head or crevices if a threat is detected [gonzalez-rivero_linking_2017, asunsolo-rivera_behaviour_2023]. We thus focus our study to estimating distances of fish to potential shelters, such as coral heads or other biological hotspots.
@@
Biologists use a metric named flight initiation distance (FID) [asunsolo-rivera_behaviour_2023] as the distance when the flight response is first triggered. From our model above, we can directly compute the FID by setting $f(x_{FID})=\alpha L_{control}$ and solving for $x_{FID}$, where $\alpha\in[0,1)$ depending on how conservative an estimate is desired. To mitigate disturbance, AUVs should attempt to operate outside the estimated FID.
@@ 2.2 Field data acquisition
Next, we placed an AprilTag next to the coral head that was used for fine-grained localization of the AUV. We used a custom AUV/ROV [girdhar_curee_2023] named CUREE, that utilizes 6 BlueROV T200 thrusters, to run several transects across the target coral head, though any AUV/ROV could be used as long as it is able to run consistent linear transects. After each pass of the AprilTag, the vehicle waited at least 1 minute in an idle state, roughly 4 m away from the AprilTag, in order to allow the fish behavior to re-acclimate. This time period was selected based on qualitative estimates from divers, but future work should study re-acclimation time periods and potential long-term acclimatization. We also recorded qualitative data that the fish did not acclimatize over the long-term, but further studies are needed to confirm this. Before the vehicle was placed in the water, we time-synchronized all the external cameras and vehicle on-board systems.

### v2

@@ 1 Introduction
Autonomous underwater vehicles (AUVs) and remotely operated vehicles (ROVs) have been proposed as scalable alternatives or supplementary systems for diver-based surveys and trawling methods to estimate species abundance, biodiversity, and animal behavior in a variety of environments ranging from coral reefs to deeper benthic communities [4, 9]. Many of these approaches rely on cameras to record videos or imagery during vehicle surveys, which can then be used to visually count animals or study behavior directly.
Recent studies of marine animal interactions with underwater vehicles [7, 8] suggest that the presence of vehicles themselves can lead to a bias in the estimation of these quantities, impacting scientists’ ability to trust those measurements. For instance, if animals are repelled or attracted to a vehicle, this will cause under- or over-estimation of their populations, respectively. However, many of these studies were conducted on animals in captivity, which often do not reflect the behaviors of animals in the wild [10].
Campbell et al. [3] provides one of the first studies of marine animal-robot interaction in the wild. They used a camera trap setup and measured fish responses to AUVs, ROVs, and Towed Vehicles (TVs). Benoit-Bird et al. [2] followed with a study about deep sea fish response to light on AUVs and ROVs by using echosounders for measurements. In both studies, their initial findings suggest that different vehicles do produce noticeable shifts in fish counts. However, due to the logistical challenges of each setup, it is difficult to reproduce and validate a variety of other hypotheses in other types of environments or species. The goals of this work is to:
@@
    Provide a theoretical framework to identify and systematically minimize measurement bias of AUVs for collecting data about marine animal behavior, that is agnostic to sensor and emittance types, which allows us to methodically test across them.
@@
    Provide and validate a simple protocol for collecting data in-situ to inform the model.
@@ 2.1 Model of animal response
As in [3], we hypothesize that certain species of fish show behavioral responses to vehicles, but that it is highly correlated with their sensing modalities. However, many of these quantities are difficult to measure and so, regardless of modality, there is some threshold distance at which their sensing or reaction is limited. It is known that many prey fish will shelter in the coral head or crevices if a threat is detected [6, 1]. We thus focus our study to estimating distances of fish to potential shelters, such as coral heads or other biological hotspots.
@@
Biologists use a metric named flight initiation distance (FID) [1] as the distance when the flight response is first triggered. From our model above, we can directly compute the FID by setting $f(x_{FID})=\alpha L_{control}$ and solving for $x_{FID}$, where $\alpha\in[0,1)$ depending on how conservative an estimate is desired. To mitigate disturbance, AUVs should attempt to operate outside the estimated FID.
@@ 2.2 Field data acquisition
Next, we placed an AprilTag next to the coral head that was used for fine-grained localization of the AUV. We used a custom AUV/ROV [5] named CUREE, that utilizes 6 BlueROV T200 thrusters, to run several transects across the target coral head, though any AUV/ROV could be used as long as it is able to run consistent linear transects. After each pass of the AprilTag, the vehicle waited at least 1 minute in an idle state, roughly 4 m away from the AprilTag, in order to allow the fish behavior to re-acclimate. This time period was selected based on qualitative estimates from divers, but future work should study re-acclimation time periods and potential long-term acclimatization. We also recorded qualitative data that the fish did not acclimatize over the long-term, but further studies are needed to confirm this. Before the vehicle was placed in the water, we time-synchronized all the external cameras and vehicle on-board systems.